OpenStudy (anonymous):

write 0.375 as an unreduced fraction then use words to name the number

OpenStudy (anonymous):

How do I do this

OpenStudy (mayankdevnani):

 very easy method: if you convert decimal into fraction.first you see that how many digits are there after decimal and just divide that number(without decimal) by the no. of digits as 0... for eg. 3.75=375/100 ,, 3.7=37/10 and so on!!! got it @Mookie?$